### v1

@@ Abstract
Fifteen years of the Fermi Large Area Telescope (LAT) data in the halo region of the Milky Way (MW) are analyzed to search for gamma rays from dark matter annihilation. Gamma-ray maps within the region of interest (, ) are modeled using point sources, the GALPROP models of cosmic-ray interactions, isotropic background, and templates of Loop I and the Fermi bubbles, and then the presence of a halo-like component is further examined. A statistically significant halo-like excess is found with a sharp peak around 20 GeV, while its flux is consistent with zero below 2 GeV and above 200 GeV. Examination of the fit residual maps indicates that a spherically symmetric halo component fits the map data well. The radial profile agrees with annihilation by the smooth NFW density profile, and may be slightly shallower than this, especially in the central region. Various systematic uncertainties are investigated, but the 20 GeV peak remains significant. In particular, the halo excess with a similar spectrum is detected even relative to the LAT standard background model, which does not depend on GALPROP or other model templates. The halo excess can be fitted by the annihilation spectrum with a mass 0.5–0.8 TeV and annihilation cross section (5–8) for the channel. This cross section is larger than the upper limits from dwarf galaxies and the canonical thermal relic value, but considering various uncertainties, especially the density profile of the MW halo, the dark matter interpretation of the 20 GeV “Fermi halo” remains feasible. The prospects for verification through future observations are briefly discussed.
@@ 1 Introduction
Even though nearly a century has passed since Fritz Zwicky inferred the existence of dark matter from the motion of galaxies in the Coma cluster [1], dark matter remains one of the greatest mysteries in astrophysics, cosmology, and fundamental physics [2, 3, 4]. Weakly interacting massive particles (WIMPs) in the mass range of 1 GeV to 100 TeV have long been considered one of the leading candidates, because they interact only through the weak interaction, and because the thermal relic amount with the annihilation cross section expected from particle physics is close to the observed dark matter density [3, 4, 5, 6]. The existence of WIMPs can be probed with direct searches by underground experiments, indirect searches using annihilation products (e.g., gamma-rays and neutrinos), and by particle accelerator experiments.
Regarding indirect searches by gamma-rays, the sensitivities of the Fermi Large Area Telescope (Fermi LAT) and ground-based air Cherenkov gamma-ray telescopes have reached a level where the intensity expected from the canonical annihilation cross section can be detected. When WIMPs directly annihilate into two photons, the line gamma-ray emission is easy to distinguish, but the branching ratio into line emission is highly uncertain. In the more mainstream annihilation mode, the Standard Model particles such as or are first generated, followed by a cascade that produces gamma rays with a continuum energy spectrum. The spectrum still shows a broad peak around a specific energy related to the WIMP mass, which contrasts with power-law spectra of many astrophysical objects. Numerous explorations have been conducted toward various astronomical objects or regions in search of such signals, including the Milky Way (MW) Galactic center (GC) [7, 8, 9], dwarf galaxies [10, 11, 12, 13] and dark subhalos [14, 15] in the MW halo, the Andromeda galaxy [16, 17], galaxy groups and clusters [18, 19, 20, 21], and the extragalactic background radiation [22, 23, 24, 25].
@@
However, the GC GeV excess can also be explained by astrophysical phenomena rather than dark matter, such as millisecond pulsars [8]. Then the true dark matter signal may appear in other photon energy ranges. The dark matter interpretation of the GC excess suggests that the radial density profile near the GC is with 1.1-1.3, which is more centrally concentrated than the standard Navarro-Frenk-White (NFW) [38] profile ( in the central region). If the true density profile is the NFW or less centrally concentrated, dark matter signals may be more easily detected in the halo region than at the GC. In 2012, the Fermi LAT team analyzed the MW halo region in search of dark matter using two years of the LAT data, which is much smaller than the amount accumulated to date. From that time until now, understanding and analysis methods for diffuse radiation in the halo region, such as the Fermi bubbles and Loop I, have also matured. For these reasons, it would be worthwhile to investigate the presence or absence of a halo-like radiation in addition to the known components based on the latest LAT data. This study reports the results of a dark matter search using 15 years of the LAT data.
@@ 2.2 Likelihood analyses
In many studies fitting a model map to the LAT data, the photon counts and the model expectations are compared in each pixel. However, the halo-like component is expected to be much smoother than the pixel scale of 0.125∘. Therefore, we divide the ROI into cells with a width of and calculate the likelihood based on photon counts and expectations in these cells. This not only speeds up calculations, but also prevents the likelihood from being unnecessarily biased by pixel-scale fine discrepancies between observed counts and model predictions. The maximum likelihood fitting is performed independently for each photon energy bin. Let be the observed photon count in the -th cell and be the number of photons expected from the model, and then the likelihood function based on Poisson statistics is
@@ 2.3 Model map templates
Since the extended sources are not many, circular regions with a radius twice that of the semimajor axis of the catalog are masked and excluded from the analysis.
@@
This parameter set is used for the inverse Compton scattering (ICS) component in the Galactic interstellar emission model (GIEM, gll_iem_v07.fits)555https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html [46, 47] recommended by the LAT team for normal point source analysis.
@@
The GALPROP model emission can be further divided into more components, and the fitting results with such a larger number of sub-components or different model parameters will be discussed in section 3.4.3.
@@
Loop I is a giant structure spanning more than 100∘ in the radio sky [48], and although it is unclear whether it is nearby supernova remnants or an outflow from the GC [49, 50], it has also been detected by LAT [51]. We adopt the geometric model of Loop I, which assumes a uniform emissivity within two shells with radii of 50–100 pc, based on the nearby hypothesis. The parameters of the geometric model are taken from the LAT team papers [37, 7], which differ slightly from the values in Wollben’s original paper [49]. The normalization factors for each of the two shells are treated as independent fit parameters ().
@@
To calculate the halo component model map, we first assume that the density profile of the MW dark matter halo is NFW, and use the parameters from the Via Lactea II simulation [52, 53]. Therefore, the density profile within the virial radius = 402 kpc is , where , 21 kpc, and . The distance from the GC to the solar system is set at 8 kpc. Annihilation gamma-ray diffuse flux is proportional to the line-of-sight integration of the gamma-ray emissivity per unit volume. If dark matter is distributed smoothly, the emissivity obeys , and the map template in this case is denoted as NFW-. According to numerical simulations, fine substructures and subhalos exist within halos composed of cold dark matter, and subhalos with high internal densities may dominate annihilation gamma-ray luminosity. If the number density of subhalos is proportional to the smooth density profile, the gamma-ray emissivity would be . This emissivity also applies when dark matter particles decay rather than annihilate. Therefore, the map template of (NFW-) will also be tested. Subhalos may be destroyed by strong tidal forces near the center of the parent halo [52], in which case emissivity in the central region would be suppressed, but this is not tested in this study. Finally, based on previous studies indicating that the density profile of () is preferred for the GC GeV excess, the model template of (NFW-), which becomes the same emissivity profile as the GC excess in the central region using the original NFW, is included in the fit for this component.
@@
The map morphologies of the point sources and GALPROP have photon energy dependencies determined by the models, while other model maps do not depend on photon energy. The initial values of in the MCMC chain are the same as the original normalization for the models of point sources and GALPROP. The initial of the isotropic component is set to , which is typical of the reported fluxes of the extragalactic gamma-ray background radiation [54]. For other components, the initial values of the MCMC chain are set to zero. During the MCMC fitting, the fitting parameters are limited to the range for known components such as point sources, the GALPROP models, the isotropic component, and Loop I, but is allowed for the halo components, which are explored as unknown components.
@@ 3.1 Construction of the Fermi bubble template
Unlike Loop I and the MW halo components, there are no parameterized analytical models for the Fermi bubbles. Therefore, the map template for the bubbles used in this study is made here, mostly following the methods adopted in the LAT team’s papers [37, 7]. As the first step, a flat (i.e., isotropic) template within the rough area of the Fermi bubbles is assumed, and it is fit to the data together with the other model components. Here, the GC GeV excess is included in the fit as a known component, but other halo components are not considered. At this stage, the fit is performed in the whole ROI, including the Galactic disk region (), to ensure an accurate fit to the GC excess. Looking at the best-fit image of the bubbles (i.e., sum of the flat template and fit residuals, smoothed by the Gaussian with ), the energy-independent boundary of the flat template is iteratively improved.
@@
The image of the Fermi bubbles obtained in this way is shown in figure 1, together with the final boundary of the flat template, in the two energy bins of 1.5 and 4.3 GeV, which are in the typical photon energy range where the Fermi bubbles are prominent. Since the template of the Fermi bubbles will only be used at in subsequent halo component searches, the Galactic disk region () is not shown. Residuals of bright point sources are visible in some places, and these are likely due to uncertainties in the catalog spectral models and PSF. The fit in this study is performed on a cell-by-cell basis with the cell size of , and PSF errors should have little effect on the results.
@@
The best-fit spectra (calculated by the best-fit in each energy bin) for all the model components by this fit are shown in figure 2. Here, it should be noted that the spectral flux shown in this figure excludes the disk region for comparison with the later analyses, though the model fitting is performed including the disk region. The spectra of the model templates are generally consistent with the results of similar analyses by the LAT team [7], indicating that there are no serious problems or errors in this analysis. In particular, the peak photon energy and flux of the GC GeV excess ( at 1∘ from the GC at 2.5 GeV) are also consistent with those reported in the literature [8].
@@
The morphology of the Fermi bubbles is also consistent with what has been reported in the literature. However, in the northern region ( 10–15∘), negative residuals are clearly visible, especially at the lower energy bin of 1.5 GeV. This feature corresponds to the regions of particularly strong flux in the gas component map of GALPROP (see Figure 13 in section 3.3 later), and is likely due to discrepancies between the GALPROP model and the data. A similar feature is seen in the LAT team’s analysis using GALPROP (Fig. 22 of ref. [37]), and the LAT team used the method of spectral component analysis to produce a bubble template without this feature. Since the purpose of the present study is not to create an accurate template of the Fermi bubbles, the following approach will be taken to explore an additional halo component.
@@
Among the two energy bins, the 4.3 GeV map is chosen as the source to create the template, because the Fermi bubbles are more clearly seen. In the map shown in Figure 1, the positive flux regions are dominated by the Fermi bubbles. Therefore, the map of positive regions, regardless of whether it is inside or outside the boundaries of the flat template, will be used as the energy-independent template for the bubbles. On the other hand, in order to eliminate the effects of discrepancies between GALPROP and the data, the map of the negative flux regions is introduced into the fit as another template. Since these two probably have different energy spectra, they are included as two independent templates.
@@
3.2 Fitting results with a halo-like component
From here, the fits will be performed excluding the Galactic disk region () to search for a halo-like component. The flat template of the Fermi bubbles is replaced with the positive and negative residual templates obtained in the previous section. The fit parameter is limited to for the positive residual template (the Fermi bubbles), but not for the negative template, because the inverted negative template (i.e., positive flux) may be favored for a better fit, if the true morphology of the Fermi bubbles is relatively flat and the discrepancy between GALPROP and the data decreases in higher energy bins,
@@ 3.2 Fitting results with a halo-like component
Next, the fit results with the GC GeV excess component (NFW-) added are shown in figure 4. The GC excess flux is the same as the previous fit that includes the Galactic plane (figure 2) at 4.3 GeV. This is how it should be, because the two residual templates are created at this energy bin. However, in higher photon energy regions, the GC excess component rises sharply and peaks at around 20 GeV, where the flux is more than 10 times higher than the previous fit. This means that extrapolating this fit to the GC would significantly exceed the observed data. These results indicate that a halo-like component exists around 20 GeV, with a density profile shallower than that assumed for the GC GeV excess, and is prominent in the region.
Therefore, to test a flatter density profile, we perform fits using the NFW- and NFW- halo models instead of the GC excess, and the results are shown in figures 5 and 6, respectively. Only one of the three halo models is used in a fit to avoid degeneracy between the similar halo maps. Since the best-fit halo flux becomes negative at the lowest photon energy bin, figure 7 shows the spectra of the three halo models on a linear plot. The flatter the density profile, the higher the mean background flux in the ROI, but in all cases, the intensity of the halo components rises sharply with photon energy between 1 GeV and 20 GeV, reaching a statistically significant (13–19) peak at the 21 GeV bin. At higher energy bins, the halo flux decreases rapidly, becoming almost zero at 200 GeV or higher.
The halo components are not completely zero at the 4.3 GeV bin, where the positive and negative templates are created from the residual map, because the GC excess component was included in the fit that yielded the residual map. The halo component flux decreases toward lower photon energies, becoming negative at 1.5 GeV. This may be attributed to the difference between the GALPROP gas component and the data, which becomes more pronounced at lower energies. Another possibility is that part of the halo component at 4.3 GeV has been absorbed in the residual map templates. If the halo component weakens toward lower energy below 4.3 GeV, the fit of the halo component could be negative to cancel out the halo component absorbed in the residual map templates.
Figure 8 illustrates the likelihood (relative to the no-halo fit) of the fits with three different halo models, to compare the goodness of fit. In the 2.5 GeV bin, the halo components are fitted to be nearly zero (figure 7), and a deep dip is seen in the figure. Around the peak photon energy of the halo excess (10–20 GeV), the flatter NFW- model fits better than NFW- at 95% confidence level (CL). On the other hand, the shallowest NFW- model fits even worse than the NFW- model (at a significance much higher than 95% CL), suggesting that a profile close to NFW- best fits the data in the region of .

### v2

@@ Abstract
Fifteen years of the Fermi Large Area Telescope (LAT) data in the halo region of the Milky Way (MW) are analyzed to search for gamma rays from dark matter annihilation. Gamma-ray maps within the region of interest (, ) are modeled using point sources, the GALPROP models of cosmic-ray interactions, isotropic background, and templates of Loop I and the Fermi bubbles, and then the presence of a halo-like component is further examined. A statistically significant halo-like excess is found with a spectral peak around 20 GeV, while its flux is consistent with zero below 2 GeV and above 200 GeV. Examination of the fit residual maps indicates that a spherically symmetric halo component fits the map data well. The radial profile agrees with annihilation by the smooth NFW density profile, and may be slightly shallower than this, especially in the central region. Various systematic uncertainties are investigated, but the 20 GeV peak remains significant. In particular, the halo excess with a similar spectrum is detected even relative to the LAT standard background model, which contains non-template patches adjusted to match the observed map. The halo excess spectrum can be fitted by annihilation with a particle mass 0.5–0.8 TeV and cross section (5–8) for the channel. This cross section is larger than the upper limits from dwarf galaxies and the canonical thermal relic value, but considering various uncertainties, especially the density profile of the MW halo, the dark matter interpretation of the 20 GeV “Fermi halo” remains feasible. The prospects for verification through future observations are briefly discussed.
@@ 1 Introduction
Even though nearly a century has passed since Fritz Zwicky inferred the existence of dark matter from the motion of galaxies in the Coma cluster [1], dark matter remains one of the greatest mysteries in astrophysics, cosmology, and fundamental physics [2, 3, 4]. Weakly interacting massive particles (WIMPs) in a mass range from 1 GeV to 100 TeV near the electroweak scale have long been considered one of the leading candidates, because such particles are naturally predicted in some particle physics theories beyond the Standard Model, and because the thermal relic amount predicted from the annihilation cross section expected for WIMPs is close to the observed dark matter density [3, 4, 5, 6]. The existence of WIMPs can be probed with direct searches by underground experiments, indirect searches using annihilation products (e.g., gamma-rays and neutrinos), and by particle accelerator experiments.
Regarding indirect searches by gamma rays, the sensitivities of the Fermi Large Area Telescope (Fermi LAT) and ground-based air Cherenkov gamma-ray telescopes have reached a level where the intensity expected from the canonical annihilation cross section can be detected. When WIMPs directly annihilate into two photons, line (monoenergetic) gamma-ray emission is easy to distinguish, but the branching ratio into line emission is highly uncertain. In the more mainstream annihilation mode, the Standard Model particles such as or are first generated, followed by a cascade that produces gamma rays with a continuum energy spectrum. The spectrum still shows a broad peak around a specific gamma-ray energy related to the WIMP mass, which contrasts with power-law spectra of many astrophysical objects. Numerous explorations have been conducted toward various astronomical objects or regions in search of such signals, including the Milky Way (MW) Galactic center (GC) [7, 8, 9], the Large and Small Magellanic Clouds (LMC and SMC) [10, 11], dwarf galaxies [12, 13, 14, 15] and dark subhalos [16, 17] in the MW halo, the Andromeda galaxy [18, 19], galaxy groups and clusters [20, 21, 22, 23], and the extragalactic background radiation [24, 25, 26, 27].
@@
However, the GC GeV excess can also be explained by astrophysical phenomena rather than dark matter, such as millisecond pulsars [8]. Then the true dark matter signal may appear in other photon energy ranges. To explain the GC excess with dark matter annihilation, the radial density profile near the GC must be with 1.1–1.3, which is more centrally concentrated than the standard Navarro-Frenk-White (NFW) [40] profile ( in the central region). If the true density profile is NFW or less centrally concentrated, dark matter signals may be more easily detected in the halo region than at the GC. In 2012, the Fermi LAT team analyzed the MW halo region in search of dark matter using two years of the LAT data [28], which is much smaller than the amount accumulated to date. From that time until now, understanding and analysis methods for diffuse radiation in the halo region, such as the Fermi bubbles and Loop I, have also matured. For these reasons, it would be worthwhile to investigate the presence or absence of a halo-like radiation in addition to the known components based on the latest LAT data. This study reports the results of a dark matter search using 15 years of the LAT data.
@@ 2.2 Likelihood analyses
The maximum likelihood fitting is performed independently for each photon energy bin. In many studies that fit a model map to the LAT data, the photon counts and the model expectations are compared in each pixel. Since the halo-like component is expected to be much smoother than the pixel scale of 0.125∘, calculating the likelihood at a coarser resolution allows computation time to be reduced without compromising dark matter search. Therefore, we divide the ROI into cells with a width of and calculate the likelihood based on photon counts and expectations in these cells. Some components of the Galactic diffuse emission, such as those from interstellar gas, possess pixel-scale fine structures, which may help separate them from others by pixel-scale likelihood calculation. On the other hand, smoothing at coarse resolutions may reduce the impact of pixel-scale mismatches between a model map and data on the likelihood estimate.
Let be the observed photon count in the -th cell and be the number of photons expected from the model, and then the likelihood function based on Poisson statistics is
@@ 2.3 Model map templates
Since the extended sources are not many, circular regions with a radius twice the semi-major axis of the catalog are masked and excluded from the analysis.
@@
This parameter set is used for the inverse Compton scattering (ICS) component in the Galactic interstellar emission model (GIEM, gll_iem_v07.fits)555https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html [48, 49], which is recommended by the LAT team for normal point source analysis.
@@
The flux ratio between pion decay and bremsstrahlung is fixed to the original model value, making the gas component a single template.
@@
The GALPROP model emission can be further divided into different components, and the fitting results with a larger number of sub-components or different model parameters will be discussed in section 3.4.4.
@@
Loop I is a giant structure spanning more than 100∘ in the radio sky [50], and although it is unclear whether it is nearby supernova remnants or an outflow from the GC [51, 52], it has also been detected by LAT [53]. We adopt the geometric model of Loop I, which assumes a uniform emissivity within two shells with radii of 50–100 pc, based on the nearby hypothesis. The parameters of the geometric model are taken from the LAT team papers [39, 7], which differ slightly from the values in Wollben’s original paper [51]. The normalization factors for each of the two shells are treated as two independent fit parameters ().
@@
To calculate the halo component model map, we first assume that the density profile of the MW dark matter halo is NFW, and use the parameters from the Via Lactea II simulation [54, 55]. Therefore, the density profile within the virial radius = 402 kpc is , where , 21 kpc, and . The distance from the GC to the solar system is set at 8 kpc, and hence . Annihilation gamma-ray diffuse flux is proportional to the line-of-sight integration of the gamma-ray emissivity per unit volume. If dark matter is distributed smoothly, the emissivity obeys , and the map template in this case is denoted as NFW-. According to numerical simulations, fine substructures and subhalos exist within halos composed of cold dark matter, and subhalos with high internal densities may dominate annihilation gamma-ray luminosity. If the number density of subhalos is proportional to the smooth density profile, the gamma-ray emissivity would be . This emissivity also applies when dark matter particles decay rather than annihilate. Therefore, the map template of (NFW-) will also be tested. Subhalos may be destroyed by strong tidal forces near the center of the parent halo [54], in which case emissivity in the central region would be suppressed. While a model for this case is not tested in this work, this possibility should be kept in mind. Previous studies on the GC GeV excess found that the density profile of () is preferred, and hence the template of with the original NFW (NFW-) is adopted as the model of the GC excess, whose emissivity becomes in the central region ().
@@
The map morphologies of the point sources and GALPROP have photon energy dependencies determined by the models, while other model maps do not depend on photon energy. The initial values of in the MCMC chain are the same as the original normalization for the models of point sources and GALPROP. The initial of the isotropic component is set to , which is typical of the reported fluxes of the unresolved extragalactic gamma-ray background radiation [56]. For other components, the initial values of the MCMC chain are set to zero. During the MCMC fitting, the fitting parameters are limited to the range for known components such as point sources, the GALPROP models, the isotropic component, and Loop I, but is allowed for the halo components, which are explored as unknown components.
@@ 3.1 Construction of the Fermi bubble template
Unlike Loop I and the MW halo components, there are no parameterized analytical models for the Fermi bubbles. Therefore, the map template for the bubbles used in this study is made here, mostly following the methods adopted in the LAT team’s papers [39, 7]. Two photon energy bins, 1.5 and 4.3 GeV, are selected for this procedure, which are in the typical photon energy range where the Fermi bubbles are prominent. As the first step, a flat (i.e., isotropic) template with sharp edges within a rough area of the Fermi bubbles is assumed, and it is fit to the data together with the other model components. Here, the GC GeV excess is included in the fit as a known component, but other halo components are not considered. At this stage, the fit is performed in the whole ROI, including the Galactic disk region (), to ensure an accurate fit to the GC excess.
@@
The best-fit image of the bubble is created from the sum of the flat template and fit residuals; the photon count map of this is converted into physical flux at each pixel using the exposure data, and then smoothed with a Gaussian filter with . The energy-independent boundary of the flat template is iteratively improved by looking at these best-fit images. The image of the Fermi bubbles obtained in this way in the two energy bins is shown in figure 1, together with the final boundary of the flat template. Since the template of the Fermi bubbles will only be used at in subsequent halo component searches, the Galactic disk region () is not shown. Residuals of bright point sources are visible in some places, and these are likely due to uncertainties in the catalog spectral models and PSF. The fit in this study is performed on a cell-by-cell basis with the cell size of , and PSF errors should have little effect on the results.
@@
The best-fit spectra of all the model components by the fit with the final flat bubble template are presented in figure 2. Here, the mean background flux within the ROI, including the Galactic disk region, is shown. For comparison with later analysis, figure 3 shows the flux excluding the disk, though the model fitting is the same as figure 2, i.e., including the disk. The spectra of the model templates are generally consistent with those of similar analyses by the LAT team [7], indicating that there are no serious problems or errors in this analysis. In particular, the peak photon energy and the flux of the GC GeV excess ( at 1∘ from the GC at 2.5 GeV) are also consistent with those reported in the literature [8].
@@
The morphology of the Fermi bubbles is also consistent with what has been reported in the literature. However, in the northern region ( 10–15∘), negative residuals are clearly visible, especially at the lower energy bin of 1.5 GeV. This feature corresponds to the regions of particularly strong flux in the gas component map of GALPROP (see figure 14 in section 3.3 later), and is likely due to discrepancies between the GALPROP model and the data. A similar feature is seen in the LAT team’s analysis using GALPROP (figure 22 of ref. [39]), and the LAT team used the method of spectral component analysis to produce a bubble template without this feature. Since the purpose of the present study is not to create an accurate template of the Fermi bubbles, the following approach will be taken to explore an additional halo component.
@@
Among the two energy bins, the 4.3 GeV map is chosen as the source to create the structured bubble template, because the Fermi bubbles are more clearly seen. In the map shown in figure 1, the positive flux regions are dominated by the Fermi bubbles. Therefore, the map of positive regions, regardless of whether it is inside or outside the boundaries of the flat template, will be used as the energy-independent template for the bubbles. On the other hand, to minimize the impact of the discrepancy between GALPROP and the data, the map of the negative flux regions is introduced into the fit as another template. Since these two probably have different energy spectra, they are included as two independent templates. Hereafter, they will be referred to as the positive and negative residual templates, though these strictly include not only the fit residual but also the flat bubble template.
@@
3.2 Fitting with a halo-like component
From here, the fits will be performed excluding the Galactic disk region () to search for a halo-like component. The flat template of the Fermi bubbles is replaced with the positive and negative residual templates obtained in the previous section. The fit parameter is limited to for the positive residual template (the Fermi bubbles), but not for the negative template, because the inverted negative template (i.e., positive flux) may be favored for a better fit, if the true morphology of the Fermi bubbles is relatively flat and the discrepancy between GALPROP and the data decreases in higher energy bins.
@@ 3.2 Fitting with a halo-like component
Next, the fit results with the GC GeV excess component (NFW-) added are shown in figure 5. The GC excess flux is the same as the previous fit that includes the Galactic plane (figure 3) at 4.3 GeV. This is how it should be, because the two residual templates are created at this energy bin. However, in higher photon energy regions, the GC excess component rises sharply and peaks at around 20 GeV, where the flux is more than 10 times higher than the previous fit. This means that extrapolating this fit to the GC would significantly exceed the observed data. These results imply that a halo-like component exists around 20 GeV and is prominent in the region, with a density profile shallower than that assumed for the GC GeV excess.
Therefore, to test a flatter density profile, we perform fits using the NFW- and NFW- halo models instead of the GC excess, and the results are shown in figures 6 and 7, respectively. Only one of the three halo models is used in a fit to avoid degeneracy between the similar halo maps. Since the best-fit halo flux becomes negative at the lowest photon energy bin, figure 8 shows the spectra of the three halo models on a linear plot. The flatter the density profile, the higher the mean background flux in the ROI, but in all cases, the intensity of the halo components rises sharply with photon energy between 1 GeV and 20 GeV, reaching a statistically significant (13–19) peak at the 21 GeV bin. At higher energy bins, the halo flux decreases rapidly, becoming almost zero at 200 GeV or higher.
The halo components are not completely zero at the 4.3 GeV bin, where the positive and negative templates are created from the residual map, because the GC excess component was included in the fit that yielded the residual map. The halo component flux decreases toward lower photon energies, becoming negative at 1.5 GeV. This may be attributed to the difference between the GALPROP gas component and the data, which becomes more pronounced at lower energies. Another possibility is that part of the halo component at 4.3 GeV has been absorbed in the residual map templates. If the halo component weakens toward lower energy below 4.3 GeV, the fit of the halo component could be negative to cancel out the part absorbed in the residual map templates.
Figure 9 illustrates the likelihood (relative to the no-halo fit) of the fits with three different halo models, to compare the goodness of fit. In the 2.5 GeV bin, the halo components are fitted to be nearly zero (figure 8), and a deep dip is seen in the figure. Around the peak photon energy of the halo excess (10–20 GeV), the NFW- model fits better than the steepest NFW- at 95% confidence level (CL). On the other hand, the shallowest NFW- model fits even worse than the NFW- (at a significance much higher than 95% CL), suggesting that a profile close to NFW- best fits the data in the region of .